OpenStudy (anonymous):

Write the sum or difference in simplest form. Jack has 4/8 of a pizza left from a party. If he eats another 1/8, how much is left? A 3/8 B 2/3 C 5/8 D 1/2

OpenStudy (anonymous):

so c

OpenStudy (anonymous):

Add the fractions together: \[\frac{ 4 }{ 8} + \frac{ 1 }{ 8 } = ?\] Then subtract that answer from 8/8: \[8/8 - \frac{ ? }{ 8 } = youranswer\]

OpenStudy (anonymous):

No, it wouldn't be c.

OpenStudy (anonymous):

Do you know the answer, or do you need more help?:)

OpenStudy (anonymous):

I thought it was c so what is it if it isn't c

OpenStudy (anonymous):

The question is asking you how much pizza will be left after you eat 4/8 + 1/8.. so: \[4/8 + 1/8 = 5/8\] You need to know how much IS LEFT, not how much YOU ATE. So: \[8/8 - 5/8 = 3/8\]